OpenStudy (anonymous):

The original price of a digital keyboard was $119.50. The keyboard was on sale for $109.94. What was the discount rate? there was no multiple choice on it and i will give a medal if answered correct

OpenStudy (anonymous):

original-sale / original *100

OpenStudy (anonymous):

119.50-109.94 / 119.50 *100 gives you 8%

OpenStudy (jdoe0001):

we know the original price was $119.50 the discounted price is $109.94 119.50-109.94 = 9.56 so. of 119.50 is the 100%, what is 9.56 in percent ? \(\large \begin{array}{ccllll} quantity&percentage\\ \hline\\ 119.50&100\\ 109.94&x \end{array}\implies \cfrac{119.50}{109.94}=\cfrac{100}{x}\)

OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

anyhow. that gave what was 109.94 but \(\large \begin{array}{ccllll} quantity&percentage\\ \hline\\ 119.50&100\\ 9.56&x \end{array}\implies \cfrac{119.50}{9.56}=\cfrac{100}{x}\)